A box contains 13 red hats, 15 white hats, 18 blue hats, and 14 brown hats. Nina randomly chooses a hat from the box.

What is the probability that it will NOT be white? Write your answer as a decimal.

Answer by jim_thompson5910(35256)

There are 13+15+18+14 = 60 hats total. There are 60 - 15 = 45 non-white hats. So the probability is 45/60 = 3/4, which in decimal form is 0.75 and in percentage form is 75%.

So the answer they want is 0.75
--------------------------------------------------------------------------------------------------------------

If you need more help, email me at jim_thompson5910@hotmail.com

You can also visit my website: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim

--------------------------------------------------------------------------------------------------------------